OpenStudy (anonymous):

Five times the difference of twice a number and six is negative twenty. Find the number.

OpenStudy (anonymous):

n=1

OpenStudy (anonymous):

5(2n-6)=-20 10n-30=-20 10n=10 n=1

OpenStudy (anonymous):

5(2x-6)= -20 =>10x-30=-20 =>10x=-20+30=10....ie,x=1